OpenStudy (anonymous):

Why can't I get this?? Please help.. What is the solution of the system? y=-3x 3x + 2y=6

OpenStudy (amistre64):

well, they tell you y = - 3x, what do we do with that?

OpenStudy (anonymous):

solve for 2 coordinates using elimination

OpenStudy (amistre64):

substitution would be a simple process is what im thinking

OpenStudy (anonymous):

How do I even come up with an answer to this? x and y?

OpenStudy (amistre64):

well, they tell you that y = -3x, so everywhere you see a y, replace it by -3x; would be a start

OpenStudy (anonymous):

I came up with -2 and 6 but im so confused on what the heck to do

OpenStudy (amistre64):

3x + 2y=6 3x + 2(-3x)=6 the rest is like the most basic algebra to solve for x

OpenStudy (amistre64):

lets see if -2, 6 fits .. 6=-3(-2) good 3(-2) + 2(6)=6 -6 + 12 = 6 good nothing else to do .. you solved it

OpenStudy (anonymous):

Ok so using substitution we plug those in and show they are correct and thank you so much for your help I get so dang confused

OpenStudy (amistre64):

youre welcome :) im going home for now so good luck

OpenStudy (anonymous):

thanks!

OpenStudy (phi):

if you know the answer, you can test that the answer works, as amiste did up above. If you don't know the answer, you have to do algebra. y=-3x 3x + 2y=6 the first equation says y is the same thing as -3x (-3 times x) if you replace y in the 2nd equation with -3*x you get 3x +2(-3x) = 6 that means 3 * x + 2*-3*x = 6 2*-3 is -6 so we can simplify to 3x -6x = 6 3 x's take away 6 x's leaves -3 x's or -3x -3x = 6 divide both sides by -3 -3x/-3 = 6/-3 -3/-3 is 1 so 1*x = 6/-3 or x= 6/-3 6 divided by -3 is -2 x= -2